OpenStudy (anonymous):

Let X=C[0,1]. For f,g in X, define d(f,g) = sup(x 0 to 1)|f(x)-g(x)|. Calculate the distance between f(x)=x and g(x) = x^2

OpenStudy (experimentx):

isn't it maximizing this function |x - x^2| in the interval 0 to 1

OpenStudy (anonymous):

not sure..my lectrer said find the max of f-g first..

OpenStudy (experimentx):

this is just a parabola x(x-1) .. vertex is at 0.5

OpenStudy (experimentx):

can you define d(f,g) again?

OpenStudy (anonymous):

d(f,g)=\[\sup _{0\le x \le 1} |f(x)-g(x)|\]

OpenStudy (experimentx):

that's definitely the supremum of |f(x) - g(x)| in the interval 0-2 ...it must be 0.5

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=plot+x+-+x^2

OpenStudy (experimentx):

* interval 0-1

OpenStudy (anonymous):

means that the distance is the max value of f-g?and also the supremum?

OpenStudy (anonymous):

It's the maximum of the absolute value. So it could be the minimum if it is further away from 0. Since 0.5 is where the max is, you plug it in to find the distance: (0.5)(1-0.5)=0.25

OpenStudy (anonymous):

oo..ok.thanks